Calculate the apparent weight of an elevator attendant (m = 84 kg) if the elevator is accelerating downward at 2.80 m/s2.


1
Expert's answer
2021-01-20T17:19:23-0500

Let's apply the Newton's Second Law of Motion:


"N-mg=-ma,""N=mg-ma=m(g-a),""N=84\ kg\cdot(9.8\ \dfrac{m}{s^2}-2.80\ \dfrac{m}{s^2})=588\ N."

Answer:

"N=588\ N."